# Probing forced responses and causal mechanisms in large-scale climate dynamics with reduced-order neural models

### v3

@@ V.3 A realistic application: radiative responses to temperature perturbations
Building on the theoretical insights introduced in the first part, the second part of the paper presented a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net radiative flux at the top-of-atmosphere (TOA). The model was trained on a long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies. The emulator focuses on the evolution of 20 principal components of the SST field and on the scalar global mean, net radiative flux; unobserved degrees of freedom are parametrized through multiplicative noise through an additional neural network. The emulator largely reproduces the stationary statistics of the system and, more importantly, allows us to reconstruct TOA radiative fluxes in two forced simulations. We inferred the direct causal links between temperature modes and TOA fluxes through a perturbation approach, akin to Green’s function-like experiments with climate models [38], targeting the so-called “pattern effect” [39]. Crucially, the analyses targeted both responses in the ensemble mean and variance. Such responses are then projected back onto the EOFs to obtain maps of sensitivity in mean and variance change of the radiative flux at the TOA to SST warming perturbations. We argued that the sensitivity map for the ensemble mean is consistent with recent studies, indicating that the emulator’s skill in reconstructing TOA fluxes under forcing arises from physically meaningful mechanisms. In contrast, the sensitivity map of the ensemble variance is a novel result, not previously examined in the literature. It highlights the potential for uncertainty quantification in perturbation experiments but it requires further investigation to better understand the underlying physical mechanisms. These findings highlight the potential of reduced-order neural emulators to perform thousands of perturbation experiments, shifting the focus of climate feedback analysis from mean responses to the responses of probability distributions.
@@
Modeling the climate system from data alone is constrained by observing only a subset of variables of a multiscale, high-dimensional dynamical system. This raises the fundamental challenge of identifying the appropriate variables (both fields and spatiotemporal scales) a fundamental problem in physics, emphasized in classical work [122] and revisited in recent studies [123]. However, large-scale climate dynamics is often well characterized by a few recurrent spatiotemporal patterns, or modes of variability. In this case, the multiscale structure of the climate system can be an advantage allowing us to derive data-driven models of the coarse-grained dynamics while treating unresolved, fast processes stochastically. Traditional reduced-order models stand out from general-purpose emulators by targeting specific climate problems from the outset, guided by prior literature, physical insight, and a clear understanding of the relevant variables and scales at play. Such data-driven modeling approach has a long history in climate science [2], though most efforts have emphasized reproducing stationary statistics or short-term forecasting [3, 4, 75] rather than forced responses. A promising perspective is to build on such established reduced-order approaches, focusing on powerful neural networks to fit deterministic and stochastic components and on tools from response theory, such as the FDT, to guide model development. Together, these strategies can enable efficient algorithms to study how full probability distributions respond to external perturbations, opening new avenues for causal inference and attribution in climate science.
@@ A.2 Causal interpretation of the response operator
Causality in natural systems is inferred through external interventions: an agent perturbs the system and the response is measured [77, 78]. Climate modeling follows the same principle through Green’s function experiments, where one variable is perturbed and the response of another is evaluated [38, 124]. Linear response theory formalizes this principle, providing a rigorous framework for causal analysis both in models and directly from data [17, 18, 114].
Because responses are inherently time-dependent, causal links through responses can be distinguished in three forms, all encoded in the time-dependent response operator: (i) direct links, (ii) indirect links, and (iii) cumulative links (or causal sensitivity). In what follows, we use the triad system in Eq. (6) as a simple example to illustrate these ideas in a clear and unambiguous way. Finally, we summarize recent approaches for estimating responses directly from data and refer to Baldovin et al. (2020) [17] for details on the theoretical foundations of this approach.
@@
Figure 8: Panel (a) Functional form of the triad system in Eq. (6). Panel (b) Graph structure of the triad system. Note from panel (b) that causality in dynamical systems can involve cycles, in contrast with the acyclic structures often assumed in applications outside physics, see e.g. [125].
Causal relations can then be probed by adding independent impulse perturbations to each degree of freedom and record the system’s response. Responses are computed up to a long time horizon $\tau_{\max}$, with discrete times $t=0,dt,2dt,\dots,\tau_{max}$. Since we are interested in causal relations, it suffices to consider the identity observable $A(x_{k}(t))=x_{k}(t)$, i.e. the response of the ensemble mean [17]. We now clarify the three different types of causal links mentioned at the start of this Section.
@@
The direct causal links are encoded in the response operator $\mathbf{R}(dt)$ evaluated at the shortest time step $t=dt$. The intuition follows directly from the graph representation in Fig. 8(b) and is rather straightforward.
@@
Causal relations in physical systems extend beyond direct interactions. A variable $x_{j}$ may affect another variable $x_{k}$ indirectly through an intermediate variable $x_{i}$, forming a pathway $x_{j}\rightarrow x_{i}\rightarrow x_{k}$. Such indirect links are central in climate dynamics [126] and in complex systems more broadly [127]. Within the response framework, these pathways are revealed by the time-dependent operator $R_{k,j}(t)$: the first non-zero response identifies the shortest causal path in the graph connecting $x_{j}$ to $x_{k}$. Crucially, if $R_{k,j}(t)=0$ for all time $t$, then no external intervention on $x_{j}$ can ever influence the behavior of $x_{k}$.
@@
Finally, Baldovin et al. [17] introduced a cumulative (in time) measure of causation between two variables, defined as $D_{k,j}=\int_{0}^{\tau_{max}}R_{k,j}(t)dt$. Such cumulative link, can be also referred to as “causal sensitivity”. Such a measure is especially relevant in practical experiments with high-dimensional models where interactions unfold across multiple spatial and temporal scales. A concrete example using climate models, is provided by the approach of Bloch-Johnson et al. (2024) [38], where causal links are inferred between patterns of surface warming and the net radiative flux at the top of the atmosphere. Their method reconstructs responses to perturbations by: (i) applying surface warming perturbations, (ii) integrating an atmospheric-only model with and without the perturbation, and (iii) computing the average (in time) difference between the two runs after removing an initial transient. In this case, the causal link emerges precisely as a cumulative (in time) atmospheric response to the surface perturbations.
We emphasize the crucial role of ensemble averages in establishing the general validity of response theory. Evaluating responses along single trajectories of nonlinear systems is strongly affected by chaos, restricting the applicability of these ideas to short times and very small perturbations. In contrast, focusing on observables, i.e. quantities averaged over the invariant density, ensures the robustness of the theory and largely removes the dependence on chaos. For a detailed discussion of this point, we refer the reader to [128]. Future efforts to estimate responses in climate science, which often rely on perturbing single trajectories, would therefore benefit from focusing on ensemble-averaged observables.
@@
Given the assumption of Markovianity it is possible to infer the system’s response operator from a long stationary trajectory using the Fluctuation-Dissipation Theorem (FDT) [17]. The FDT formula provides a link between stationary and perturbed dynamics by predicting forced responses from stationary statistics alone [129]. Specifically:
@@
where $A(x_{k}(t))$ is a general observable, $\rho(\mathbf{x})$ is the stationary distribution. $\langle\cdot\rangle$ refers to ensemble average, approximated through temporal averages. $\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}$ represents the so-called “score function” evaluated at the initial state, i.e. on the attractor [15]. Therefore, in principle the response operator can be computed from data if the score $\nabla\ln\rho(\mathbf{x})$ is available. In practice, the score is often inaccessible, and most applications rely on the quasi-Gaussian (qG) approximation, by replacing $\rho(\mathbf{x})$ with a Gaussian distribution $\rho^{G}(\mathbf{x})$ with same mean and covariances [130, 48]. For responses in the ensemble mean—which are sufficient to infer causal links—the qG approximation achieves high skill even for nonlinear systems [121, 17, 114]. Recent work [114] has demonstrated that generative modeling techniques can be used to efficiently estimate the score directly from data, leading to high skill in computation of higher-order responses (up to the 5th moment) in strongly nonlinear regimes. Other fruitful avenues lie in frameworks at the intersection of Markov State Modeling and Koopman operator theory such as [131].
It is worth emphasizing that the central limitation discussed in the main text, i.e. the need to choose the proper variables, remains relevant for the inference of response operators via FDT as well (and for any causal inference methodology); see the discussion in [17].
@@ Appendix B Relevance of the the triad system for geophysical turbulence
Here, we motivate the choice of the triad model analyzed in Section III as a building block for geophysical turbulent flows. We further present and test the associated analytical stochastic model for the slow variable $x_{1}$ (as discussed in Section III.3.1), illustrating that Markovian reduced-order models can sometimes be derived from partial observations. For detailed mathematical derivations and additional examples, we refer to the original work by Majda and collaborators [46, 60, 61, 48, 16].
@@ B.1 Link to geophysical turbulence
Projecting many geophysical turbulent flows, such as quasigeostrophic models, onto a complete set of orthogonal basis functions (spectral or empirical modes such as EOFs), give rise to the following generic form for the mode amplitudes:
@@
where $\mathbf{F}$ is an external forcing, $\mathbf{L}$ is a linear operator and $\mathbf{B}(\mathbf{u},\mathbf{u})$ is a nonlinear, quadratic operator. The linear operator $\mathbf{L}=\mathbf{L}^{\prime}+\mathbf{D}$, is formed by a skew-symmetric operator $\mathbf{L}^{\prime}$ and by a diagonal dissipation matrix $\mathbf{D}$ with negative real entries. The quadratic nonlinearity should be energy-conserving, satisfying $\mathbf{u}\cdot\mathbf{B}(\mathbf{u},\mathbf{u})=0$ [132, 60, 48]. Eq. (16) is the typical abstract form of dry cores of atmospheric models [67]. Galerkin truncations of turbulent systems often reduce to this structure, as in the Lorenz ’63 and Charney–DeVore models [132, 133, 68]. $\mathbf{u}\in\mathbb{R}^{N}$ represents the full turbulent geophysical system, with $N$ infinite in principle and finite but very large in applications ($N\gg 1$). The state $\mathbf{u}$ is then decomposed into slow and fast components, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, i.e. $\mathbf{x}$ will represent the first $n$ components of $\mathbf{u}$ and $\mathbf{y}$ the next NnN-n.
We consider the simple case of $N=3$: one large-scale resolved mode $x_{1}$ and two fast modes $\mathbf{y}=(x_{2},x_{3})$. Majda and collaborators [46] considered an exact projection of the slow mode $x_{1}$ and the two fast modes $(x_{2},x_{3})$ onto the dynamics in Eq. (16). This gives rise to three equations, defined by a forcing term, linear and quadratic self- and cross-interaction across slow and fast time scales. The triad model in Eq. (6) is then obtained by approximating the quadratic self-interactions of the fast modes as a damping and stochastic forcing, as: $\gamma_{i}\epsilon^{-1}x_{i}+\sigma_{i}\epsilon^{-1/2}\xi_{i}$, with $i=2,3$. $\xi_{i}$ is a standard Gaussian white noise process. $\epsilon$ sets the time-scale separation between the slow and fast modes. The key assumption is that the detailed nonlinear self-interactions among fast modes are not important if the goal is to capture the coarse-grained slow dynamics. This approximation transforms the original deterministic equations into a stochastic system. The specific version used in Eq. (6) is further simplified by setting to zero a few interactions, such as the linear cross- and self-interactions of the fast modes as well as the external forcing terms.
Therefore, the relevance of the triad system (Eq. (6)) to this study is that it provides a minimal multiscale setting with linear and quadratic nonlinear interaction of the kind that naturally arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs. For additional details see also Section 4 of [16].
@@ B.2 Analytical stochastic model for
Starting from the triad model in Eq. (6), Majda and collaborators [60] showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ using standard projection techniques [48, 46, 47]. The mode reduction predicts the functional form of the dynamics governing the resolved, slow variable [60, 61, 62, 63]. The resulting stochastic equation is
@@
where the coefficients are expressed in terms of parameters of the triad system in Eq. (6) as: $a=L_{11}+\epsilon(\frac{I^{2}\sigma_{2}^{2}}{2\gamma^{2}_{2}}-\frac{L_{12}^{2}}{\gamma_{2}}-\frac{L_{13}^{2}}{\gamma_{3}})$; $b=-2\epsilon\frac{L_{12}I}{\gamma_{2}}$; $c=\epsilon\frac{I^{2}}{\gamma_{2}}$; $B=-I\epsilon^{1/2}\frac{\sigma_{2}}{\gamma_{2}}$; $L=-B\frac{L_{12}}{I}$; $\tilde{F}=-\frac{AB}{2}$; $\sigma=\epsilon^{1/2}L_{13}\frac{\sigma_{3}}{\gamma_{3}}$. $W_{2}$ and $W_{3}$ are independent Wiener processes. Interestingly, the equation for the coarse-grained variables $x_{1}$ display a structure richer than that of the full triad system in Eq. (6). In addition to the original linear and quadratic interactions, the reduced equation includes new terms: an effective forcing $\tilde{F}$, a cubic nonlinearity, and both additive and multiplicative noise. The additive noise results from the linear interactions across fast modes, while the multiplicative noise reflects the advection of the slow mode by the fast modes [67]. Thus, Eq. (17) demonstrates that multiplicative noise arises from theoretical considerations, rather than being introduced ad hoc, providing a firm basis for its use in reduced-order stochastic climate models. The scalar model above has served as a prototype in several contexts: it has been used to test the Fluctuation–Dissipation Theorem in climate dynamics [81, 114] as well as for inverse modeling of the leading principal component of a quasi-geostrophic model and the observed variability associated with the North Atlantic Oscillation (NAO), demonstrating its relevance for theoretical studies in climate dynamics.
@@
We evaluate the model in Eq. (17) as done in Section III, so to reproduce the stationary statistics of the slow variable $x_{1}$ as well as its response to perturbations as in the original triad system in Eq. (6). The results are reported in Figure 9. The analysis show that, given $\epsilon\ll 1$, it is possible to build a Markovian model of the slow variable $x_{1}$ reproducing both stationary and perturbed statistics as in the original system.
Finally, as shown in this Section, rigorous model reduction strategies such as those developed by Majda and collaborators [61] and by Roberts [134, 135] (among others [2]) offer invaluable theoretical guidelines for multiscale climate modeling. While such strategies where traiditionally developed in the case when the equation of motions are known, they can possibly serve as inspiration and leading to new directions in future work involving data-driven emulators.
Finally, as shown in this Section, rigorous model reduction frameworks, such as those developed by Majda and collaborators [61] and by Roberts [134, 135], among others [2], provide invaluable theoretical guidance for multiscale climate modeling. While these strategies were traditionally developed for systems with known equations of motion, they may also inspire future work aimed at incorporating such principles into data-driven-only emulators.

### v4

@@ V.3 A realistic application: radiative responses to temperature perturbations
Building on the theoretical insights introduced in the first part, the second part of the paper presented a stochastic neural emulator for the coarse-grained dynamics of sea surface temperature (SST) and net radiative flux at the top-of-atmosphere (TOA). The model was trained on a long unforced simulations of a coupled climate model and designed with a specific physical task in mind: capturing radiative feedbacks to SST anomalies. The emulator focuses on the evolution of 20 principal components of the SST field and on the scalar global mean, net radiative flux; unobserved degrees of freedom are parametrized as multiplicative noise through an additional neural network. The emulator largely reproduces the stationary statistics of the system and allows us to reconstruct TOA radiative fluxes in two forced simulations. We inferred the direct causal links between temperature modes and TOA fluxes through a perturbation approach, akin to Green’s function-like experiments with climate models [blochjohnsonetal2024], targeting the so-called “pattern effect” [Dong2019AttributingPacific]. Such responses are then projected back onto the EOFs to obtain maps of sensitivity in mean and variance change of the radiative flux at the TOA to SST warming perturbations. We discussed similarities of the sensitivity map for the ensemble mean with recent studies, indicating that the emulator’s skill in reconstructing TOA fluxes under forcing arises from physically meaningful mechanisms.We also identified limitations. In particular, the variance response is sensitive to the training of the multiplicative-noise component, restricting the emulator’s reliability in perturbation experiments beyond those targeting the ensemble mean. We outline possible paths forward to resolve this issue. These findings highlight the potential of reduced-order neural emulators to perform thousands of perturbation experiments, eventually shifting the focus of climate feedback analysis from mean responses to the responses of probability distributions.
VI Conclusion
@@
Modeling the climate system from data alone is constrained by observing only a subset of variables of a multiscale, high-dimensional dynamical system. This raises the fundamental challenge of identifying the proper variables (both fields and spatiotemporal scales), a fundamental problem in physics emphasized in classical work [Onsager] and revisited in recent studies [Lucente]. Such choices are not unique and are largely determined by the study’s scientific goal rather than by the pursuit of an unattainable perfect model [Smith2001]. In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is dominated by a few recurrent spatiotemporal patterns, or modes of variability, and their interactions. Thus, at least for certain tasks, the multiscale structure of the climate system can be an advantage, allowing us to focus on the coarse-grained dynamics while treating unresolved, fast processes stochastically. Traditional reduced-order models stand out from general-purpose emulators by targeting specific climate problems from the outset, guided by prior literature, physical insight, and a clear understanding of the relevant variables and scales at play. Such data-driven modeling approach has a long history in climate science [LucariniChekroun], though most efforts have emphasized reproducing stationary statistics or short-term forecasting [Penland89, Penland95, KONDRAENSO] rather than forced responses. A promising perspective is to build on such established reduced-order approaches, including those developed by Majda and Qi [majdaSIAM, QiMajdaChaos], while leveraging flexible neural networks architectures to discover low-dimensional latent representations [latentPierre, PierreLatent] and fit deterministic and stochastic components. Furthermore, tools from response theory such as the FDT [FabCausal, LudovicoFabriAndre, ValerioData2] can be used to guide model development and eventually lead to reduced-order models that efficiently capture responses in the mean and variance to external perturbations, advancing causal inference and attribution analyses in climate science.
@@ A.2 Causal interpretation of the response operator
Causality in natural systems is inferred through external interventions: an agent perturbs the system and the response is measured [IsmaelNew, Aurell]. Climate modeling follows the same principle through Green’s function experiments, where one variable is perturbed and the response of another is evaluated [blochjohnsonetal2024, LucariniPRL]. Linear response theory formalizes this principle, providing a rigorous framework for causal analysis both in models and directly from data [Baldovin, FabCausal, LudovicoFabriAndre].
Because responses are inherently time-dependent, causal links through responses can be distinguished in three forms, all encoded in the time-dependent response operator: (i) direct links, (ii) indirect links, and (iii) cumulative links (or causal sensitivity). In what follows, we use the triad system in Eq. (6) as a simple example to illustrate these ideas in a clear and unambiguous way. Finally, we summarize recent approaches for estimating responses directly from data and refer to Baldovin et al. (2020) [Baldovin] for details on the theoretical foundations of this approach.
@@
Figure 8: Panel (a) Functional form of the triad system in Eq. (6). Panel (b) Graph structure of the triad system. Note from panel (b) that causality in dynamical systems can involve cycles, in contrast with the acyclic structures often assumed in applications outside physics, see e.g. [Kusner].
Causal relations can then be probed by adding independent impulse perturbations to each degree of freedom and record the system’s response. Responses are computed up to a long time horizon $\tau_{\max}$, with discrete times $t=0,dt,2dt,\dots,\tau_{max}$. Since we are interested in causal relations, it suffices to consider the identity observable $A(x_{k}(t))=x_{k}(t)$, i.e. the response of the ensemble mean [Baldovin]. We now clarify the three different types of causal links mentioned at the start of this Section.
@@
The direct causal links are encoded in the response operator $\mathbf{R}(dt)$ evaluated at the shortest time step $t=dt$. The intuition follows directly from the graph representation in Figure 8(b) and is rather straightforward.
@@
Causal relations in physical systems extend beyond direct interactions. A variable $x_{j}$ may affect another variable $x_{k}$ indirectly through an intermediate variable $x_{i}$, forming a path $x_{j}\rightarrow x_{i}\rightarrow x_{k}$. Such indirect links are central in climate dynamics [RungeCausalEffects] and in complex systems more broadly [Ay]. Within the response framework, these paths are revealed by the time-dependent operator $R_{k,j}(t)$: the first non-zero response identifies the shortest causal path in the graph connecting $x_{j}$ to $x_{k}$. Crucially, if $R_{k,j}(t)=0$ for all time $t$, then no external intervention on $x_{j}$ can ever influence the behavior of $x_{k}$.
@@
Finally, Baldovin et al. [Baldovin] introduced a cumulative (in time) measure of causation between two variables, defined as $D_{k,j}=\int_{0}^{\tau_{max}}R_{k,j}(t)dt$. Such cumulative link, can be also referred to as “causal sensitivity”. Such a measure is especially relevant in practical experiments with high-dimensional models where interactions unfold across multiple spatial and temporal scales. A concrete example using climate models, is provided by the approach of Bloch-Johnson et al. (2024) [blochjohnsonetal2024], where causal links are inferred between patterns of surface warming and the net radiative flux at the top of the atmosphere. Their method reconstructs responses to perturbations by: (i) applying surface warming perturbations, (ii) integrating an atmospheric-only model with and without the perturbation, and (iii) computing the average (in time) difference between the two runs after removing an initial transient. In this case, the causal link emerges precisely as a cumulative (in time) atmospheric response to the surface perturbations.
We emphasize the crucial role of ensemble averages in establishing the general validity of response theory. Evaluating responses along single trajectories of nonlinear systems is strongly affected by chaos, restricting the applicability of these ideas to short times and very small perturbations. In contrast, focusing on observables, i.e. quantities averaged over the invariant density, ensures the robustness of the theory and largely removes the dependence on chaos. For a detailed discussion of this point, we refer the reader to [FALCIONI1995481]. Future efforts to estimate responses in climate science, which often rely on perturbing single trajectories, would therefore benefit from focusing on ensemble-averaged observables.
@@
Given the assumption of Markovianity it is possible to infer the system’s response operator from a long stationary trajectory using the Fluctuation-Dissipation Theorem (FDT) [Baldovin]. The FDT formula provides a link between stationary and perturbed dynamics by predicting forced responses from stationary statistics alone [Falcioni]. Specifically:
@@
where $A(x_{k}(t))$ is a general observable, $\rho(\mathbf{x})$ is the stationary distribution. $\langle\cdot\rangle$ refers to ensemble average, approximated through temporal averages. $\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}$ represents the so-called “score function” evaluated at the initial state, i.e. on the attractor. Therefore, in principle the response operator can be computed from data if the score $\nabla\ln\rho(\mathbf{x})$ is available. In practice, the score is often inaccessible, and most applications rely on the quasi-Gaussian (qG) approximation, by replacing $\rho(\mathbf{x})$ with a Gaussian distribution $\rho^{G}(\mathbf{x})$ with same mean and covariances [Leith, MajdaBook]. For responses in the ensemble mean the qG approximation achieves high skill even for nonlinear systems [GRITSUN, Baldovin, LudovicoFabriAndre]. Recent work [LudovicoFabriAndre] has demonstrated that generative modeling techniques can be used to efficiently estimate the score directly from data, leading to high skill in computation of higher-order responses in strongly nonlinear regimes. Other interesting new proposals have been recently formulated in the context of Koopman theory and can be used for future exploration of forced responses from data alone [ValerioData1, ValerioData2, FDTKoopman].
It is worth emphasizing that the central limitation discussed in the main text, i.e. the need to choose the proper variables for the system to be Markovian, remains relevant for the inference of response operators via FDT as well (and for any causal inference methodology); see the discussion in [Baldovin].
@@ Appendix B Relevance of the the triad system for geophysical turbulence
Here, we motivate the choice of the triad model analyzed in Section III as a building block for geophysical turbulent flows. We further present and test the associated analytical stochastic model for the slow variable $x_{1}$ (as discussed in Section III.3.1). For detailed mathematical derivations and additional examples, we refer to the original work by Majda and collaborators [NormalForms, MTV1, MTV2, MajdaBook, majdaSIAM].
@@ B.1 Link to geophysical turbulence
Projecting many geophysical turbulent flows, such as quasigeostrophic models, onto a complete set of orthogonal basis functions (spectral or empirical modes such as EOFs), give rise to the following generic form for the modes amplitudes:
@@
where $\mathbf{F}$ is an external forcing, $\mathbf{L}$ is a linear operator and $\mathbf{B}(\mathbf{u},\mathbf{u})$ is a nonlinear, quadratic operator. The linear operator $\mathbf{L}=\mathbf{L}^{\prime}+\mathbf{D}$, is formed by a skew-symmetric operator $\mathbf{L}^{\prime}$ and by a diagonal dissipation matrix $\mathbf{D}$ with negative real entries. The quadratic nonlinearity should be energy-conserving, satisfying $\mathbf{u}\cdot\mathbf{B}(\mathbf{u},\mathbf{u})=0$ [Lorenz, MTV1, MajdaBook]. Eq. (16) is the typical abstract form of dry cores of atmospheric models [MajdaPerspective]. Galerkin truncations of turbulent systems often reduce to this structure, as in the Lorenz ’63 and Charney–DeVore models [Lorenz, Charney, CROMMELIN]. $\mathbf{u}\in\mathbb{R}^{N}$ represents the full turbulent geophysical system, with $N$ infinite in principle and finite but very large in applications ($N\gg 1$). The state $\mathbf{u}$ is then decomposed into slow and fast components, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, i.e. $\mathbf{x}$ represents the first $n$ components of $\mathbf{u}$ and $\mathbf{y}$ the next NnN-n.
We consider the simple case of $N=3$: one large-scale resolved mode $x_{1}$ and two fast modes $\mathbf{y}=(x_{2},x_{3})$. We then project the three modes onto the dynamics in Eq. (16). This gives rise to three equations, defined by a forcing term, linear and quadratic self- and cross-interaction across slow and fast time scales. The triad model in Eq. (6) is then obtained by approximating the quadratic self-interactions of the fast modes as a damping and stochastic forcing, as: $\gamma_{i}\epsilon^{-1}x_{i}+\sigma_{i}\epsilon^{-1/2}\xi_{i}$, with $i=2,3$. $\xi_{i}$ is a standard Gaussian white noise process. $\epsilon$ sets the time-scale separation between the slow and fast modes. The key assumption is that the detailed nonlinear self-interactions among fast modes are not important if the goal is to capture the coarse-grained slow dynamics. This approximation transforms the original deterministic equations into a stochastic system. The specific version used in Eq. (6) is further simplified by setting to zero a few interactions, such as the linear cross- and self-interactions of the fast modes as well as the external forcing terms.
Therefore, the relevance of the triad system (Eq. (6)) to this study is that it provides a minimal multiscale setting with linear and quadratic nonlinear interaction of the kind that naturally arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs. Furthermore, the model is characterized by fast decays of correlations, and generates PDFs that are regular and quasi-Gaussian, as often seen in the analysis of large-scale modes of climate variability [MajdaStructuralStability]. For additional details see [NormalForms] and Section 4 of [majdaSIAM].
@@ B.2 Analytical stochastic model for
Starting from the triad model in Eq. (6), Majda and collaborators [MTV1] showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ using standard projection techniques [MajdaBook, NormalForms, Majda2010]. The mode reduction predicts the functional form of the dynamics governing the resolved, slow variable [MTV1, MTV2, MTV3, MTV4]. The resulting stochastic equation is:
@@
where the coefficients are expressed in terms of parameters of the triad system in Eq. (6) as: $a=L_{11}+\epsilon(\frac{I^{2}\sigma_{2}^{2}}{2\gamma^{2}_{2}}-\frac{L_{12}^{2}}{\gamma_{2}}-\frac{L_{13}^{2}}{\gamma_{3}})$; $b=-2\epsilon\frac{L_{12}I}{\gamma_{2}}$; $c=\epsilon\frac{I^{2}}{\gamma_{2}}$; $B=-I\epsilon^{1/2}\frac{\sigma_{2}}{\gamma_{2}}$; $L=-B\frac{L_{12}}{I}$; $\tilde{F}=-\frac{AB}{2}$; $\sigma=\epsilon^{1/2}L_{13}\frac{\sigma_{3}}{\gamma_{3}}$. $W_{2}$ and $W_{3}$ are independent Wiener processes. Interestingly, the equation for the coarse-grained variable $x_{1}$ displays a structure richer than that of the full triad system in Eq. (6). In addition to the original linear and quadratic interactions, the reduced equation includes new terms: an effective forcing $\tilde{F}$, a cubic nonlinearity, and both additive and multiplicative noise. The additive noise results from the linear interactions across fast modes, while the multiplicative noise reflects the advection of the slow mode by the fast modes [MajdaPerspective]. Thus, Eq. (17) demonstrates that multiplicative noise arises from theoretical considerations, rather than being introduced ad hoc, providing a firm basis for its use in reduced-order stochastic climate models. The scalar model above has been used to test the Fluctuation–Dissipation Theorem in climate dynamics [MajdaStructuralStability, LudovicoFabriAndre] as well as for inverse modeling of the leading principal component of a quasi-geostrophic model and the observed variability associated with the North Atlantic Oscillation (NAO), demonstrating its relevance for theoretical studies in climate dynamics.
@@
We evaluate the model in Eq. (17) as done in Section III, so to reproduce the stationary statistics of the slow variable $x_{1}$ as well as its response to perturbations as in the original triad system in Eq. (6). The results are reported in Figure 9. The analysis shows that, given $\epsilon\ll 1$, it is indeed possible to build a Markovian model of the slow variable $x_{1}$ reproducing both stationary and perturbed statistics as in the original system.
Finally, as shown in this Section, rigorous model reduction frameworks, such as those developed by Majda and collaborators [MTV2] and by Roberts [Roberts2006, ROBERTS200812], among others [LucariniChekroun], provide invaluable theoretical guidance for multiscale climate modeling. While these strategies were traditionally developed for systems with known equations of motion, they may also inspire future work aimed at incorporating such principles into data-driven-only emulators.